OpenStudy (anonymous):

Daniela has $35.10 in dimes and quarters. She has 1 more dimes than quarters. How many quarters does she have?

OpenStudy (kuoministers):

we need to write 2 equations to solve this

OpenStudy (kuoministers):

let dimes = y let quarters = x

OpenStudy (anonymous):

(.10)y+.25x=35.10 but that doesn't make sense if she has more dimes than quarters

OpenStudy (kuoministers):

(x+1)10 + (25x) = 3510

OpenStudy (anonymous):

O ok.

OpenStudy (kuoministers):

that would be the first equation right?

OpenStudy (anonymous):

Yes that makes sense.

OpenStudy (kuoministers):

because she has 1 more dime than quarters

OpenStudy (kuoministers):

(x+1)10 + (25x) = 3510 10x+10 + 25x =3510 35x+10=3510 35x=3500 can you work this out now?

OpenStudy (kuoministers):

you there? :)

OpenStudy (anonymous):

100 right?

OpenStudy (anonymous):

I got a little confused at first and I went back and I had made a small mistake.

OpenStudy (kuoministers):

ahh yes 100 is the answer